OpenStudy (anonymous):

a 2.0 kg cart moving due east at 6.0 m/s collides with a 3.0 kg cart moving due west. the cart sticks togther and come to rest afterthe collision. what was the inital speed of the 3.0 kg cart?

OpenStudy (anonymous):

m1=2kg,v1=6m/s . m2=3kg v2=? m1*v1+m2*v2=(m1+m2)V V=0 answer should be negative ,because it goes to opposite direction.

OpenStudy (anonymous):

Apply Law of Conservation of Momentum Take the direction of east = + \[M_{1}U_{1} + M_{2}U_{2} = M_{1}V_{1} +M_{2}V_{2}\] \[2 \times 6 - 3 \times U_{1} = (M_{1}+M_{2})V\] since it is at rest, V = 0 \[12 = 3U_{1}\] \[U_{1} = 4ms\]